A project generates $100,000 cash flow per year (similar to a dividend when looking at stocks). The investor thinks the end of year one cash flow will equal $100,000 times 1.03. The investor thinks these cash flows may grow at 3% per year. The investor wants to earn a 13% interest rate on this investment. Compute the possible apartment building value today. [constant growth model CH8].
Answer #1

Cash Flow Generate d in Year 1 = D1 = 100000

Growth rate of cash flow = g = 3%

Rate of Return = r = 13%

According to Gordon's Growth model,

Value today = D1/(r - g) = 100000 / (0.13 - 0.03) = $1000000